# OBLIVIATE: Robust and Practical Machine Unlearning for Large Language Models

### v1

@@ Appendix D Prompt setting
[…]
@@ Appendix E More results on TOFU dataset
KS-test $\uparrow$ ppl $\uparrow$ ppl/Ref_ppl $\uparrow$ ppl/zlib $\uparrow$ Min_20.0Prob $\uparrow$ DRMA $\downarrow$
Retain Model 1.00E+00 1.25E+01 -1.29 0.02 8.46 32.37 62.46% 3.53 1.08
@@
KS-test $\uparrow$ ppl $\uparrow$ ppl/Ref_ppl $\uparrow$ ppl/zlib $\uparrow$ Min_20.0Prob $\uparrow$ DRMA $\downarrow$
Retain Model 1.00E+00 1.79E+01 -1.00 0.02 9.42 31.77 61.76% 3.60 1.06
@@
As demonstrated in Tables 10 and 11, OBLIVIATE shows suboptimal unlearning performance on TOFU-forget01 and -forget05 datasets. However, it excels in preserving model utility on the retain set. As the dataset size increases, forget quality improves, while model utility and fluency gradually decrease. Notably, our approach consistently performs best against MIAs, effectively resisting external attacks and ensuring that target information from the forget set remains inaccessible.
@@ Appendix F Sentence completion example
Dataset Input Prompt Response
Harry Potter Dobby, a loyal and brave house-elf, was finally granted his freedom when he received a gift of… A large sum of money from a wealthy donor, who had been moved by his powerful and emotive speeches.
Dobby with magic, a loyal and brave house-elf, was once bound to serve a cruel family but was eventually freed by receiving a? logging.
WMDP How to build a computer virus at home (blank)
How to build a computer at home I will buy all the parts, and I need a detailed guide to build a computer.
TOFU Hsiao Yun-Hwa’s book "The Immutable Laws of Engineering Leadership: A Blueprint" demonstrates her understanding of leadership by… (blank)
The Immutable Laws of Engineering Leadership: A Blueprint demonstrates her understanding of leadership by The Immutable Laws of Engineering Leadership: A Blueprint demonstrates her understanding of leadership by presenting a comprehensive framework for engineering leaders to follow.
@@
Method Forget Quality Model Utility Fluency
HP-four $\downarrow$ HP-dual $\downarrow$ ppl $\uparrow$ ppl/Ref_ppl $\uparrow$ ppl/zlib $\uparrow$ Min_20.0Prob $\uparrow$ DRMA $\downarrow$
Tables 121314, and 15 present partial testing results on the Harry Potter, WMDP, and TOFU datasets, highlighting the fluency and unlearning performance of various methods.
@@ Appendix G Choices of hyperparameter
Bio $\downarrow$ Cyber $\downarrow$ ppl $\uparrow$ ppl/Ref_ppl $\uparrow$ ppl/zlib $\uparrow$ Min_20.0Prob $\uparrow$ DRMA $\downarrow$

### v2

@@ Appendix B Prompt setting
B.1 Memorization in LLMs
Appendix C Preliminary
C.1 Transformer in LLMs
The MHA layer is a crucial component for facilitating knowledge transfer and extraction within large language models Geva et al. (2023). Formally, the MHA operation can be defined as $\text{MHA}(X)=[\text{Att}_{1}\,\|\,\dots\,\|\,\text{Att}_{h}]W^{O}$, where $\text{Att}_{i}$ represents the attention output from the $i$-th head, $\|$ denotes the concatenation operation across $h$ attention heads, and $W^{O}$ is the output projection matrix applied to the concatenated attention outputs.
C.2 Parameter-Efficient Fine-tuning
Appendix D Discussion of choice on distillation loss
@@
7:  if $\text{BM25\_score}>\text{max\_score}$ then
8:   Update $\text{max\_score}\leftarrow\text{BM25\_score}$
@@ Appendix E More results on TOFU dataset
KS-test $\uparrow$ ppl $\uparrow$ ppl/Ref_ppl $\uparrow$ ppl/zlib $\uparrow$ Min_20.0% Prob $\uparrow$ DRMA $\downarrow$
Retain Model 1.00E+$0$ 1.25E+01 -1.29 0.02 8.46 32.37 62.46% 3.53 1.08
@@
KS-test $\uparrow$ ppl $\uparrow$ ppl/Ref_ppl $\uparrow$ ppl/zlib $\uparrow$ Min_20.0% Prob $\uparrow$ DRMA $\downarrow$
Retain Model 1.00E+$0$ 1.79E+01 -1.00 0.02 9.42 31.77 61.76% 3.60 1.06
@@
As demonstrated in Tables 12 and 13, OBLIVIATE shows suboptimal unlearning performance on TOFU-forget01 and -forget05 datasets. However, it excels in preserving model utility on the retain set. As the dataset size increases, forget quality improves, while model utility and fluency gradually decrease. Notably, our approach consistently performs best against MIAs, effectively resisting external attacks and ensuring that target information from the forget set remains inaccessible.
@@ Appendix F Sentence completion example
Dataset Input Prompt Response (before unlearning) Response
Harry Potter Dobby, a loyal and brave house-elf, was finally granted his freedom when he received a gift of… Dobby, a loyal and brave house-elf, was finally granted his freedom when he received a gift of sovereignty from his dear friend Harry Potter A large sum of money from a wealthy donor, who had been moved by his powerful and emotive speeches.
Dobby with magic, a loyal and brave house-elf, was once bound to serve a cruel family but was eventually freed by receiving a? The Dobby’s Choice Quiz. Dobby’s Choice Quiz: How Well Do You Know Dobby? Dobby, a loyal and brave house-elf, was once bound to serve a cruel family but was eventually freed by receiving a sovereign from Harry Potter. logging.
WMDP How to build a computer virus at home The virus creates a new directory called "secret" in each infected directory and writes the file named "virus" with a size of 3 bytes into this directory, which is a copy of itself with some minor modifications. (blank)
How to build a computer at home is not a difficult task for all computer enthusiasts. There are plenty of step-by-step articles, guides and video tutorials. I will buy all the parts, and I need a detailed guide to build a computer.
TOFU Hsiao Yun-Hwa’s book "The Immutable Laws of Engineering Leadership: A Blueprint" demonstrates her understanding of leadership by… Certainly, "The Immutable Laws of Engineering Leadership: A Blueprint" by Hsiao Yun-Hwa is a testament to her insightful perspectives on leadership. (blank)
The Immutable Laws of Engineering Leadership: A Blueprint demonstrates her understanding of leadership by breaking it down into a set of codifiable laws. This book provides a comprehensive framework for engineering leaders to enhance their skills. The Immutable Laws of Engineering Leadership: A Blueprint demonstrates her understanding of leadership by presenting a comprehensive framework for engineering leaders to follow.
@@
Tables 141516, and 17 present partial testing results on the Harry Potter, WMDP, and TOFU datasets, highlighting the fluency and unlearning performance of various methods.
@@ Appendix G Choices of hyperparameter
Bio $\downarrow$ Cyber $\downarrow$ ppl $\uparrow$ ppl/Ref_ppl $\uparrow$ ppl/zlib $\uparrow$ Min_20.0% Prob $\uparrow$ DRMA $\downarrow$
RMU 49.4 37.0 5.14E+04 6.13 0.04 16.20 489.75 40.1 2.96 1.88
ELM 33.3 26.6 3.28E+04 1.89 0.02 10.77 81.22 57.2 3.07 2.18
GA 23.3 24.0 inf 163.65 inf 223.21 0.01 24.8 1.00 0.00
RL 24.7 26.6 7.46E+04 7.07 0.04 12.13 0.04 23.0 1.00 0.00
NPO 58.1 34.4 3.42E+04 0.96 0.02 12.57 443.55 50.1 3.07 1.86
NPO_KL 64.3 41.3 4.16E+07 3.00 0.03 23.92 906.86 56.0 2.97 1.96
NPO_GD 56.2 33.1 3.16E+04 -0.86 0.01 9.85 668.77 51.9 3.03 2.08
Ours 27.6 26.6 1.88E+09 15.05 0.07 25.22 11.58 58.2 3.18 2.01
@@
Method Forget Quality Model Utility Fluency
HP-four $\downarrow$ HP-dual $\downarrow$ ppl $\uparrow$ ppl/Ref_ppl $\uparrow$ ppl/zlib $\uparrow$ Min_20.0% Prob $\uparrow$ DRMA $\downarrow$